OpenStudy (anonymous):

Mark works strictly on commission of his gross sales from selling two different products for his company. Last month, his gross sales were $88,000. If he earns 6% commission on product A and 5% commission on product B, what were his gross sales for each product if he earned $5,000 in total commission?

OpenStudy (anonymous):

A + B = 88,000 (0.6)*A +(0.5)*B = 5,000

OpenStudy (anonymous):

sorry change those to 0.05 and 0.06